# Short-term plasticity recalls forgotten memories through a trampoline mechanism

### v1

@@ Abstract
We analyze continuous Hopfield associative memories augmented by additional, rapid short-term associative synaptic plasticity. Through the cavity method, we determine the boundary between the retrieval and forgetting, or spin-glass phase, of the network as a function of the fraction of stored memories and the neuronal gain. We find that short-term synaptic plasticity yields marginal improvements in critical memory capacity. However, through dynamical mean field theory, backed by extensive numerical simulations, we find that short-term synaptic plasticity has a dramatic impact on memory retrieval above the critical capacity. When short-term synaptic plasticity is turned on, the combined neuronal and synaptic dynamics descends a high-dimensional energy landscape over both neurons and synapses. The energy landscape over neurons alone is thus dynamic, and is lowered in the vicinity of recent neuronal patterns visited by the network, just like the surface of a trampoline is lowered in the vicinity of regions recently visited by a heavy ball. This trampoline-like reactivity of the neuronal energy landscape to short-term plasticity in synapses can lead to the recall of stored memories that would otherwise have been forgotten. This occurs because the dynamics without short-term plasticity transiently moves towards a stored memory before departing away from it. Thus short-term plasticity, operating during the transient, lowers the energy in the vicinity of the stored memory, eventually trapping the combined neuronal and synaptic dynamics at a fixed point close to the stored memory. In this manner, short-term plasticity enables the recall of memories that would otherwise be forgotten, by trapping transients that would otherwise escape. We furthermore find an optimal time constant for short-term synaptic plasticity, matched to the transient dynamics timescale, to empower recall of forgotten memories.
@@ I Introduction
At its core, a neural network is made up by a set of computational units, the neurons, and the connection mediating the influence of a neuron’s activity onto another one, the synapses. Given these ingredients, two paradigmatic lines of research have been explored across the years. A first line of inquiry focuses on the types of computation and the nature of the dynamics of the neurons for a given synaptic structure [gao2015simplicity, vyas2020computation], which is held fixed throughout the deployment of the neurons’ dynamics. The seminal works on associative memories [little1974existence, hopfield1982, hopfield1984, amit1985, amit1987statistical, gardner1988optimal] and their dense versions [abbott1987storage, gardner1987multiconnected, horn1988capacities] belong to this line of inquiry. Another approach consists instead in considering synapses and neurons as a composite dynamical system, where both elements coevolve on similar timescales, with the dynamics of the synapses influencing the dynamics of the neurons, and vice versa. In biological networks, the existence of short-term synaptic plasticity driven by pre-synaptic activity is well documented [zucker2002short, abbott2004synaptic], but an expanding body of work demonstrates that also a form of short-time associative synaptic plasticity, depending on both pre- and post- synaptic neural activity is possible [malenka1991postsynaptic, volianskis2013different]. This form of short-term synaptic plasticity is thought to play a pivotal role in the development of working memory [lansner2023fast]. For artificial neural networks, recent work has showed that allowing fluctuations of the weights of a trained architecture can improve their performance [hinton1987using, miconi2018differentiable, miconi2020backpropamine, miconi2023learning], and it has been reported that a combination of astrocyte-mediated higher order interactions and short-term synaptic plasticity improves the retrieval capabilities of biologically plausible associative memories [kozachkov2025neuron]. Understanding limits and possibilities enabled by coupled neural-synaptic computation is thus a much desirable goal to improve both our modeling capabilities of biological networks and the performance of artificial ones.
@@
To our knowledge, Dong and Hopfield [dong1992dynamic] were the first to consider a model recurrent neural network endowed exclusively with short-term synaptic plasticity in 1992, and showed how the latter can give rise to simple forms of connectivity structure under an external stimulus. Recent work by Clark and Abbott [clark2024dynamic] showed that, when short-term associative synaptic plasticity is added on top of a random connectivity matrix, a variety of interesting dynamical regimes can be obtained, including the possibility to freeze the otherwise chaotic dynamics of the network into a given configuration. Moreover, it has been shown that superimposing short-term associative synaptic plasticity on a random connectivity background allows, after an external time-dependent drive is applied, to retain simple dynamical patterns for relatively large times [wakhloo2025associative]. In these works, the impact of short-term synaptic plasticity on a flat, or highly unstructured, connectivity background is considered. However, both in artificial and biological neural networks, short-term synaptic plasticity may act upon a structured connectivity matrix, shaped in the past by a much slower dynamical process, such as a training algorithms for artificial neural network or long-term plastic development in the brain [morris2003long, lynch2004long]. We currently lack a framework to understand the interplay between long-term and short-term learning in neural computation.
How does short-term synaptic plasticity operate when superimposed to a fixed, structured connectivity background? This is the question addressed by the present work. As a minimal model neural network with structured connectivity background, we consider a recurrent neural network (RNN) with pairwise, fixed Hopfield couplings, to which we superimpose a dynamically evolving, plastic synaptic matrix. Hopfield networks constitute a celebrated model for the implementation of associative memories [little1974existence, hopfield1982, hopfield1984, ramsauer2020hopfield, krotov2023new]. Drawing inspiration from the Hebbian learning rule [hebb2005organization], these models store in their connectivity matrix a certain number of patterns. Already in the absence of short-term synaptic plasticity, the dynamics of the neurons in the network is remarkably complicated [amit1985, amit1987statistical, Kuhn1993], and it depends on the fraction of pattern stored in the connectivity matrix with respect to the number of neurons. If this fraction is below a given threshold, termed the critical capacity, different patterns partition the phase space of the network into distinct attractors. In this retrieval phase, robust recovery of a given pattern is thus possible at long times, provided that the network shares some degree of alignment (measured by the so-called overlap) with the target pattern in its initial state. As the fraction of patterns exceeds the critical capacity, different basins of attraction interfere with each other, giving rise to spurious states. Even when initialized in the vicinity of a target pattern, the network is driven toward these spurious states, which bear no overlap with the target pattern. In this catastrophic forgetting phase, spurious basins organize into a hierarchical structure, typical of spin-glass models. The dynamics of the network in this phase is rather subtle: even though at long times the overlap with the target pattern decays to zero, the dynamical overlap exhibits a nonmonotonic behavior at intermediate times, [clark2025transient]. This transient retrieval is a reflection of the complicated structure of the basins of attraction of the network.
In this work, we study the impact of short term, associative synaptic plasticity on a Hopfield-RNN. We employ a combination of static cavity method, dynamical mean field theory and extensive numerical simulations to understand the impact of synaptic plasticity on the retrieval capabilities of the network. Our salient results are summarized in Fig. 1.
@@
Figure 1: Retrieval capabilities of Hopfield recurrent neural network with short-term associative synaptic plasticity. The heatmap shows the long-time overlap of the neuron’s output with a target pattern, $m^{*}_{\infty}$, when the network is initialized in its vicinity. The long-time overlap is measured using dynamical mean-field theory, for different values of the fraction of stored pattern $\alpha$, and the short-term plasticity strength $k$. The red dashed line denotes the phase boundary, $\alpha_{c}(k)$, between the retrieval phase of the network and a forgetting phase, as obtained from a static analysis of the network behavior. When $\alpha>\alpha_{c}(k)$, without synaptic plasticity or for moderate values of $k$, the long time overlap of the network with the target pattern decays to zero. This region is where catastrophic forgetting takes place (white region). However, as the short-term plasticity strength is increased, we report the occurrence of plastic retrieval: long-time recovery of the target pattern takes place even when the network is loaded above its critical capacity, $\alpha>\alpha_{c}(k)$. The neural gain is set to be $\gamma=3.4$.
@@
We find that the presence of associative synaptic plasticity effectively acts on the fixed points of the network as a self-coupling in the retrieval phase. The strength of such a self coupling is determined by the activity of the network itself, because of the plastic feedback. From a static viewpoint, such a self-coupling leads to a marginal increase to the estimated critical capacity of the neural network (the red, dashed line in Fig. 1. However, synaptic plasticity has far richer consequences when the full dynamical history of the network is taken into account. Indeed, we discover that synaptic plasticity can exploit transient memory retrieval at the early times of the dynamics to converge to a fixed point in the vicinity of a target pattern, granting robust recovery. This finding indicates that synaptic plasticity can significantly increase the size of the basins of attractions of metastable states around the memories of the network, an effect which goes unnoticed under the conventional static analysis. We christen such a dynamical, synapses-enabled recovery plastic retrieval. The combined neural-synaptic dynamics of the network descends over a high-dimensional, composite energy landscape, over both neurons and synapses. The energy landscape of the neurons alone is constantly reshaped by the synaptic activity, very much like a trampoline is constantly deformed by a ball rolling through it. The short-term plasticity feedback can thus lower the energy in the vicinity of the target pattern, eventually trapping the network to a fixed point which share a significant overlap with the desired memory. Furthermore, for a given level of synaptic plasticity strength, we find that there is an optimal timescale that maximizes the impact of plastic retrieval, suggesting the presence of a complicated interplay between the shape of the basins of attraction and the timescales of dynamical evolution of the system. Our work sheds light on the interplay between short-term and long-term memory in neural computations, and opens the way to principled applications of enabling weight fluctuations in associative memories and biologically-plausible learning.
@@ II Model and dynamics
We consider a recurrent neural network of $N$ neurons in continuous time, following the graded-response model [Kuhn1991statistical, Kuhn1993]. Each neuron $i$ is described through the time evolution of its membrane potential, $x_{i}(t)$. The latter evolves according to the following leaky-integrator equation of motion,
@@
The first term, $J_{ij}$, represents a static connectivity matrix, modeling a form of long-term memory process. In this work, we consider a long-term connectivity matrix $J_{ij}$ that implements an Hebbian learning rule [hebb1949, hopfield1982, amit1985]: $J_{ij}$ is a Wishart matrix storing $P=\alpha N$, $N$-dimensional memory patterns $\{\xi_{i}^{\mu}\}_{\mu=1}^{P}$, namely
@@
The second term on the right hand side of Eq. (3), $A_{ij}(t)$, represents a dynamic, short-term plasticity matrix, inspired by models of activity-dependent synapses [dong1992dynamic, tsodyks1997, mongillo2008, clark2024dynamic]. It evolves on a plasticity timescale $p$, according to its own Hebbian rule, tracking the current neural activity rather than a set of fixed stored patterns. The dynamics of $A_{ij}(t)$ reads
@@
The two contributions on the right hand side represent respectively a linear decay term that eventually leads past activity to be washed out from $A_{ij}(t)$, and a synaptic reinforcement term that strengthens the connections among neuron that are simultaneously firing. The plasticity strength $k$ sets the intensity of this short-term reinforcement effect. This model thus couples the neural dynamics in Eq. (1) with the synaptic dynamics in Eq. (5), allowing us to study the interplay between fixed-pattern storage and activity-dependent synaptic modulation. If we were to choose the long-term connectivity matrix $J_{ij}$ as a random matrix with independent Gaussian entries, we would recover the model studied by Clark and Abbot [clark2024dynamic]. For $J_{ij}=0$, we recover instead the model with purely short-term synaptic plasticity, studied originally by Dong and Hopfield [dong1992dynamic].
In the absence of short-term synaptic plasticity, for $k=0$, the salient features of the model in Eq. (1) are well known [Kuhn1991statistical, Kuhn1993]. The most relevant parameter describing the retrieval capabilities of the network is the overlap $m^{\mu^{*}}(t)$ with respect to a target, or condensed, pattern $\xi^{\mu^{*}}_{i}$, defined as
@@ III Retrieval phase with short-term synaptic plasticity
We start by observing that the coupled neuron-synaptic dynamics, Eqs. 1, 2, 3, 4 and 5, minimizes a Lyapunov (or energy) function $\mathcal{L}\left[\{\phi_{i}\}_{i=1}^{N},\,\{A_{ij}\}_{i,j=1}^{N}\right]$ which is defined, up to an additive constant. as
@@
Indeed, we show in App. A that
@@
and that the inequality is saturated only at the fixed points of the dynamics of the network. The first term in Eq. (7) involves the coupling energy among the different neurons’ outputs through the connectivity matrix $W_{ij}$. The second term incorporates the membrane potential leak of each neuron, and the last term is a quadratic potential that confines the short-term plastic matrix $A_{ij}$. If $J_{ij}=0$, then we recover the purely plastic case considered by Dong and Hopfield [dong1992dynamic], while in the limit $k\to 0$, we recover the energy of a recurrent neural network with purely long-term Hebbian plasticity studied by Kühn [Kuhn1991statistical, Kuhn1993]. A neural network with short-term synaptic plasticity and Hebbian learning thus evolves in a high-dimensional energy landscape over an extended set of degrees of freedom. If we considered the restriction of the energy landscape $\mathcal{L}$ to the neuron alone, we see that such a landscape is constantly reshaped by the evolution of short-term synaptic plasticity, as the dynamics unfolds. The energy of recently visited neuronal configurations is lowered by means of the evolution of the short-term plasticity matrix $A_{ij}$. This mechanism is akin to a trampoline being deformed by a ball rolling through it. As it will be seen in Sec. IV.1, such a feedback mechanism between neural and synaptic activity is crucial in affecting the retrieval capabilities of the network.
The properties of the network at the fixed point of the dynamics, Eqs. 1, 2, 3, 4 and 5, can be studied using the static cavity method [mezard1987, del2014cavity], originally introduced in the study of spin-glass systems. The derivation is detailed in App. B, and it allows to reduce the system of $N+N^{2}$ equation for the fixed points of the dynamics of the full neuronal network to a one-dimensional problem for the effective fixed point condition of a tagged neuron in the system. The effect of short-term synaptic plasticity, as well as the influence from the other neurons, can be integrated out from the fixed point equations, and their influence can be described by means of a set of order parameter which, as customary in a mean-field setting, are self-consistently determined from the properties of the population of neurons at the fixed point itself. In particular, the cavity method allows to determine the overlap $m^{*}$ of the network with a target, condensed pattern at the fixed point, as
@@
where the output $\phi$ is the output from the tagged neuron in the network at the fixed point, and satisfies the equation
@@
where $g(x)\equiv\tanh\gamma x$, and $\xi^{*}$ is a discrete random variable, sampled uniformly from the set $\{-1,1\}$, representing a mean-field description of one of the condensed patterns. The scalar $z$ is a Gaussian noise of mean zero and unit variance, which represents the cumulative influence of all the other neurons on the cavity neuron. The average $\langle\cdot\rangle_{z,\xi^{*}}$ is an average over the realizations of the Gaussian noise and the pattern $\xi^{*}$. Averages over different realizations of the noise $z$ and the target pattern $\xi^{*}$ are statistically equivalent to averages over the population of neurons at the fixed point, under different statistical realizations of the $N$-dimensional patterns $\{\xi_{i}^{\mu}\}$. The strength of the fluctuations of the noise $z$ are set by the variance $\sigma^{2}$. The third term contributing to the argument of $g$ in Eq. (10) describes instead the backpropagated response of the system to the cavity neuron. It is proportional to the neuron output, $\phi$ and it is screened by a constant $\Gamma$. The variance $\sigma^{2}$ and the screening factor $\Gamma$ read, respectively,
@@
The properties of the fixed point determined above depend on the realizations on the noise $z$. The statistics of this noise is taken to be Gaussian, but this assumption can break down due to the development of a nonlinear instability of the network against small random perturbations to the neurons’ membrane potential. This nonlinear instability signals the transition to a spin-glass phase [amit1985, amit1987statistical], and it is studied in App. B.1. There, we derive a condition for the validity of the cavity calculation, which reads

### v2

@@ Abstract
We analyze continuous Hopfield associative memories augmented by additional, rapid short-term associative synaptic plasticity. Through the cavity method, we determine the boundary between the retrieval and forgetting, or spin-glass, phase of the network as a function of the fraction of stored memories and the neuronal gain. We find that short-term synaptic plasticity yields marginal improvements in critical memory capacity. However, through dynamical mean field theory, backed by extensive numerical simulations, we find that short-term synaptic plasticity has a dramatic impact on memory retrieval above the critical capacity. When short-term synaptic plasticity is turned on, the combined neuronal and synaptic dynamics descend a high-dimensional energy landscape over both neurons and synapses. The energy landscape over neurons alone is thus dynamic, and is lowered in the vicinity of neuronal patterns recently visited by the network, just like the surface of a trampoline is lowered in the vicinity of regions recently visited by a heavy ball. This trampoline-like reactivity of the neuronal energy landscape to short-term plasticity in synapses can lead to the recall of stored memories that would otherwise have been forgotten. This occurs because the dynamics without short-term plasticity transiently moves toward a stored memory before departing from it. Thus, short-term plasticity, operating during the transient, lowers the energy in the vicinity of the stored memory, eventually trapping the combined neuronal and synaptic dynamics at a fixed point close to the stored memory. In this manner, short-term plasticity enables the recall of memories that would otherwise be forgotten, by trapping transients that would otherwise escape. We furthermore find an optimal time constant for short-term synaptic plasticity, matched to the transient dynamics, that empowers the recall of forgotten memories.
@@ I Introduction
At its core, a neural network is made up of a set of computational units, the neurons, and the connection mediating the influence of a neuron’s activity onto another one, the synapses. Given these ingredients, two paradigmatic lines of research have been explored across the years. A first line of inquiry focuses on the types of computation and the nature of the dynamics of the neurons for a given synaptic structure [17, 54], which is held fixed throughout the deployment of the neurons’ dynamics. The seminal works on associative memories [33, 23, 24, 3, 5, 18] and their dense versions [2, 19, 25] belong to this line of inquiry. Another approach consists instead in considering synapses and neurons as a composite dynamical system, where both elements coevolve on similar timescales, with the dynamics of the synapses influencing the dynamics of the neurons, and vice versa. In biological networks, the existence of short-term synaptic plasticity driven by pre-synaptic activity is well documented [56, 1], but an expanding body of work demonstrates that a form of short-time associative synaptic plasticity, depending on both pre- and post- synaptic neural activity, is also possible [36, 53]. This form of short-term synaptic plasticity is believed to play a crucial role in the development of working memory [32]. For artificial neural networks, recent work has shown that allowing fluctuations of the weights of a trained architecture can improve their performance [22, 41, 40, 42], and it has been reported that a combination of astrocyte-mediated higher order interactions and short-term synaptic plasticity improves the retrieval capabilities of biologically plausible associative memories [27]. Also, interestingly, such short-term plasticity has recently been demonstrated in quantum optical realizations of Hopfield associative memories [37, 38]. Understanding limits and possibilities enabled by coupled neural-synaptic computation is thus a much desirable goal for improving both our modeling capabilities of biological networks and the performance of artificial ones.
@@
To our knowledge, Dong and Hopfield [10] were the first to consider a model recurrent neural network endowed exclusively with short-term synaptic plasticity in 1992, and showed how the latter can give rise to simple forms of connectivity structure under an external stimulus. Recent work by Clark and Abbott [7] showed that, when short-term associative synaptic plasticity is added on top of a random connectivity matrix, a variety of interesting dynamical regimes can be obtained, including the possibility to freeze the otherwise chaotic dynamics of the network into a given configuration. Moreover, it has been shown that superimposing short-term associative synaptic plasticity on a random connectivity background allows, after an external time-dependent drive is applied, to retain simple dynamical patterns for relatively large times [55]. In these works, the impact of short-term synaptic plasticity on a flat, or highly unstructured, connectivity background is considered. However, in both artificial and biological neural networks, short-term synaptic plasticity may act upon a structured connectivity matrix, shaped in the past by a much slower dynamical process, such as training algorithms for artificial neural network or long-term plastic development in the brain [44, 35]. We currently lack a framework to understand the interplay between long-term and short-term learning in neural computation.
@@
Figure 1: Panel (a): schematic representation of a standard associative memory. The potentials of the neurons $x_{i}(t)$ are dynamical variables, while the synaptic weights $W_{ij}$, which contains the patterns stored by the network and mediate the interactions among different neurons, are constant. Panels (b-d): Illustration of transient retrieval and eventual catastrophic forgetting in a standard associative memory. The three snapshots show the temporal evolution of the activity pattern (black star, given by $\{\phi_{1},\,\phi_{2}\}$, with $\phi_{i}\equiv\phi(x_{i})\equiv\tanh\gamma x_{i}$) of a network made by $2$ neurons within the energy landscape $\mathcal{L}\equiv\mathcal{L}(\phi_{1},\phi_{2};W_{12})$, when initialized in the vicinity of a target pattern (red cross). The neural activity initially gets closer to the target pattern, but it eventually escapes away from the target’s neighborhood. Panel (e): schematic representation of an associative memory with short term synaptic plasticity studied in this work. In this case, both the neurons’ potentials and the synapses are dynamical variables. Panels (f-h): short-term synaptic plasticity unlocks a trampoline mechanism, which empowers pattern retrieval. The three panels shows three snapshots of the activity pattern of a network of two neurons with short-term synaptic plasticity. As discussed in Sec. III, the energy surface $\mathcal{L}\equiv\mathcal{L}(\phi_{1},\phi_{2};W_{12}(t))$, over the possible network’s activity patterns for a given instantaneous value of the evolving synaptic weights $W_{ij}(t)$, is a dynamical quantity, which depends on the current value of the time-dependent synaptic couplings $W_{ij}(t)$. This dynamical energy landscape is lowered in regions recently visited by the neuron’s activity pattern. The latter is then trapped in the vicinity of the target memory, thus enabling retrieval in a situation where the standard network is led to forgetting. Panel (i): phase diagram for the retrieval capabilities of a continuous Hopfield associative memory with short-term associative synaptic plasticity, obtained from our analysis in Sec. III and Sec. IV. The heatmap shows the long-time overlap of the neuron’s output with a target pattern, $m^{*}_{\infty}$, when the network is initialized in its vicinity. The long-time overlap is measured using dynamical mean-field theory, for different values of the fraction of stored pattern $\alpha$, and the short-term plasticity strength $k$. The red dashed line denotes the phase boundary, $\alpha_{c}(k)$, between the retrieval phase of the network and a forgetting phase, as obtained from a static analysis of the network behavior. When $\alpha>\alpha_{c}(k)$, without synaptic plasticity or for moderate values of $k$, the long time overlap of the network with the target pattern decays to zero. This region is where catastrophic forgetting takes place (white region). However, as the short-term plasticity strength is increased, we report the occurrence of plastic retrieval: long-time recovery of the target pattern takes place even when the network is loaded above its critical capacity, $\alpha>\alpha_{c}(k)$. The neural gain is set to be $\gamma=3.4$. Details on how panels (b-d), (f-h) were realized are given in App. A.
How does short-term synaptic plasticity operate when superimposed on a fixed, structured connectivity background? This is the question addressed by the present work. As a minimal model neural network with structured connectivity background, we consider a recurrent neural network (RNN) with pairwise, fixed Hopfield couplings, to which we superimpose a dynamically evolving, plastic synaptic matrix. Hopfield networks constitute a celebrated model for the implementation of associative memories [33, 23, 24, 45, 29]. Drawing inspiration from the Hebbian learning rule [21], these models store in their connectivity matrix a certain number of patterns. Already in the absence of short-term synaptic plasticity, the dynamics of the neurons in the network is remarkably complicated [3, 5, 30], and it depends on the fraction of pattern stored in the connectivity matrix with respect to the number of neurons. If this fraction is below a given threshold, termed the critical capacity, different patterns partition the phase space of the network into distinct attractors. In this retrieval phase, robust recovery of a given pattern is thus possible at long times, provided that the network shares some degree of alignment (measured by the so-called overlap) with the target pattern in its initial state. As the fraction of patterns exceeds the critical capacity, different basins of attraction interfere with each other, giving rise to spurious states. Even when initialized in the vicinity of a target pattern, the network is driven toward these spurious states, which bear no overlap with the target pattern. In this catastrophic forgetting phase, spurious basins organize into a hierarchical structure, typical of spin-glass models. The dynamics of the network in this phase is rather subtle: even though at long times the overlap with the target pattern decays to zero, the dynamical overlap exhibits a nonmonotonic behavior at intermediate times, [8]. This transient retrieval is a reflection of the complex structure of the basins of attraction of the network (see Fig. 1(a-d) for a schematic representation of the network and a visual illustration of transient retrieval in a small network composed by two neurons).
@@
In this work, we study the impact of short term, associative synaptic plasticity on a Hopfield-RNN. Our salient results are summarized in Fig. 1. We show that the combined neural-synaptic dynamics of the network descends over a high-dimensional, composite energy landscape, over both neurons and synapses. The energy landscape of the neurons alone is constantly reshaped by short-term synaptic activity, very much like a trampoline is constantly deformed by a ball rolling through it. The short-term plasticity feedback can thus lower the energy in the vicinity of the target pattern, eventually trapping the network to a fixed point which shares a significant overlap with the desired memory (such a trampoline mechanism is schematically illustrated in Fig. 1 (f-h) for a small network of two neurons). By employing a combination of static cavity method, dynamical mean field theory and extensive numerical simulations, we construct a phase diagram, displayed in Fig. 1 (i), which describes how short-term synaptic plasticity affects the retrieval capabilities of the network. We find that, when the network has a single fixed point in the vicinity of each pattern (the so-called retrieval phase), the presence of short-term associative synaptic plasticity effectively acts as a self-coupling, whose strength is determined by the activity of the network itself, because of the plastic feedback. From a static point of view, this self-coupling leads to a marginal increase in the estimated critical capacity of the neural network, as shown by the red dashed line in Fig. 1 (i). However, short-term synaptic plasticity has far richer consequences when the full dynamics of the network is taken into account. Indeed, we discover that synaptic plasticity can exploit transient memory retrieval at the early times of the dynamics to converge to a fixed point in the vicinity of a target pattern, granting robust recovery. This finding indicates that synaptic plasticity can significantly increase the size of the basins of attractions of metastable states around the memories of the network, an effect which goes unnoticed under the conventional static analysis. We christen such a dynamical, synapses-enabled recovery plastic retrieval. Furthermore, for a given level of synaptic plasticity strength, we find that there is an optimal timescale that maximizes the impact of plastic retrieval, suggesting the presence of a complicated interplay between the shape of the basins of attraction and the timescales of dynamical evolution of the system. Our work sheds light on the interplay between short-term and long-term memory in neural computation, and opens the way to principled applications of enabling weight fluctuations in associative memories and biologically-plausible learning.
@@ II Model and dynamics
We consider a recurrent neural network of $N$ neurons in continuous time, following the graded-response model [31, 30]. Each neuron $i$ is described by the time evolution of its membrane potential, $x_{i}(t)$. The latter evolves according to the following leaky-integrator equation of motion,
@@
The first term, $J_{ij}$, represents a static connectivity matrix, modeling a form of long-term memory process. In this work, we consider a long-term connectivity matrix $J_{ij}$ that implements an Hebbian learning rule [20, 23, 3]: $J_{ij}$ is a Wishart matrix storing $P=\alpha N$ NN-dimensional memory patterns $\{\xi_{i}^{\mu}\}_{\mu=1}^{P}$, namely
@@
The second term on the right hand side of Eq. (3), $A_{ij}(t)$, represents a dynamic, short-term plasticity matrix, inspired by models of activity-dependent synapses [10, 52, 43, 7]. It evolves on a plasticity timescale $p$, according to its own Hebbian rule, tracking the current neural activity rather than a set of fixed stored patterns. The dynamics of $A_{ij}(t)$ reads
@@
The two contributions on the right hand side represent respectively a linear decay term that eventually leads past activity to be washed out from $A_{ij}(t)$, and a synaptic reinforcement term that strengthens the connections among neurons that are simultaneously firing. The plasticity strength $k$ sets the intensity of this short-term reinforcement effect. This model thus couples the neural dynamics in Eq. (1) with the synaptic dynamics in Eq. (5), allowing us to study the interplay between fixed-pattern storage and activity-dependent synaptic modulation. If we were to choose the long-term connectivity matrix $J_{ij}$ to be a random matrix with independent Gaussian entries, we would recover the model studied by Clark and Abbot [7]. For $J_{ij}=0$, we recover instead the model with purely short-term synaptic plasticity, studied originally by Dong and Hopfield [10].
In the absence of short-term synaptic plasticity, for $k=0$, the salient features of the model in Eq. (1) are well known [31, 30]. The most relevant parameter describing the retrieval capabilities of the network is the overlap $m^{\mu^{*}}(t)$ with respect to a target, or condensed, pattern $\xi^{\mu^{*}}_{i}$, defined as
@@ III Retrieval phase with short-term synaptic plasticity
We start by observing that the coupled neurons-synapses dynamics, Eqs. 1, 2, 3, 4 and 5, minimizes a Lyapunov (or energy) function $\mathcal{L}\left[\{\phi_{i}\}_{i=1}^{N},\,\{A_{ij}\}_{i,j=1}^{N}\right]$ which is defined, up to an additive constant, as
@@
Indeed, we show in App. B that
@@
and that the inequality is saturated only at the fixed points of the dynamics of the network. The first term in Eq. (7) involves the coupling energy among the different neurons’ outputs through the connectivity matrix $W_{ij}$. The second term incorporates the membrane potential leak of each neuron, and the last term is a quadratic potential that confines the short-term plastic matrix $A_{ij}$. If $J_{ij}=0$ (see Eq. (3)), then we recover the purely plastic case considered by Dong and Hopfield [10], while in the limit $k\to 0$, we recover the energy of a recurrent neural network with purely long-term Hebbian plasticity studied by Kühn [31, 30]. A neural network with short-term synaptic plasticity and Hebbian learning thus evolves in a high-dimensional energy landscape over an extended set of degrees of freedom. If we consider the restriction of the energy landscape $\mathcal{L}[\{\phi_{i}\}_{i=1}^{N};\,\{W_{ij}(t)\}_{i,j=1}^{N}]$ to the neurons’ activity alone, for a given instantaneous value of the synaptic weights $W_{ij}(t)$, we see that such a landscape is constantly reshaped by the evolution of short-term synaptic plasticity, as the dynamics unfolds. The energy of recently visited neuronal configurations is lowered by means of the evolution of the short-term plasticity matrix $A_{ij}$. As mentioned in the Introduction, this mechanism is akin to a trampoline being deformed by a ball rolling through it. As it will be seen in Sec. IV.1, such a feedback mechanism between neural and synaptic activity is crucial in affecting the retrieval capabilities of the network.
The properties of the network at the fixed point of the dynamics, Eqs. 1, 2, 3, 4 and 5, can be studied using the static cavity method [39, 9], originally introduced in the study of spin-glass systems. The derivation is detailed in App. C, and it allows us to reduce the system of $N+N^{2}$ equations for the fixed points of the dynamics of the full neuronal network to a one-dimensional problem for the effective fixed point condition of a tagged neuron in the system. The effect of short-term synaptic plasticity, as well as the influence from the other neurons, can be integrated out from the fixed point equations, and their influence can be described by means of a set of order parameters that, as customary in a mean-field setting, are self-consistently determined from the properties of the population of neurons at the fixed point itself. In particular, the cavity method allows to determine the overlap $m^{*}$ of the network with a target, condensed pattern at the fixed point, as
@@
where the output $\phi$ is the output of the tagged neuron in the network at the fixed point, and satisfies the equation
@@
where $g(x)\equiv\tanh\gamma x$, and $\xi^{*}$ is a discrete random variable, sampled uniformly from the set $\{-1,1\}$, representing a mean-field description of one of the condensed patterns. The scalar $z$ is a Gaussian noise of mean zero and unit variance, which represents the cumulative influence of all the other neurons on the cavity neuron. The average $\langle\cdot\rangle_{z,\xi^{*}}$ is an average over the realizations of the Gaussian noise and the pattern $\xi^{*}$. Averages over different realizations of the noise $z$ and the target pattern $\xi^{*}$ are statistically equivalent to averages over the population of neurons at the fixed point, under different statistical realizations of the $N$-dimensional patterns $\{\xi_{i}^{\mu}\}$. The strength of the fluctuations of the noise $z$ is set by the variance $\sigma^{2}$. The third term contributing to the argument of $g$ in Eq. (10) describes instead the backpropagated response of the system to the cavity neuron. It is proportional to the neuron output, $\phi$ and it is screened by a constant $\Gamma$. The variance $\sigma^{2}$ and the screening factor $\Gamma$ read, respectively,
@@
The properties of the fixed point determined above depend on the realizations on the noise $z$. The statistics of this noise is taken to be Gaussian, but this assumption can break down due to the development of a nonlinear instability of the network against small random perturbations to the neurons’ membrane potential. This nonlinear instability signals the transition to a spin-glass phase [3, 5], and it is studied in App. C.1. There, we derive a condition for the validity of the cavity calculation, which reads